### v1

# Stability of Thin Shell and Wormhole Configurations: Schwarzschild, Schwarzschild - (Anti-) de Sitter, and FLRW Spacetimes

Abstract
The stability of thin shell wormholes and black holes to linearized spherically symmetric perturbations about a static equilibrium is analyzed. Thin shell formalism is explored and junctions formed from combinations of Schwarzschild, Schwarzschild - de Sitter, and Schwarzschild - anti-de Sitter, as well as Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes are considered. The regions of stability for these different combinations are thoroughly described and plotted as a function of mass ratios of the Schwarzschild masses and radii of the wormhole throats. A taxonomy of the qualitative features of the various configurations and parameter spaces is developed, illustrating the stability regions when present. The considered wormholes are all found to be unstable in the causal region.
preprint: APS/123-QED
I Introduction
Einstein’s field equations allow for the mathematical existence of wormholes as exact solutions. A theoretical framework of constructing wormhole solutions is the thin shell or the Darmois-Israel formalism established in [1, 2] and used extensively elsewhere [3, 4, 5, 6]. This method involves the assumption that the throat of a wormhole is infinitesimally short and the energy density therein is confined to an infinitesimally thin region known as the thin shell. Using these assumptions, it is possible to derive an equation of motion for the radius of the wormhole throat using the difference in the extrinsic curvature at the throat. From this, stability conditions can be derived using an effective potential based on the equation of motion of the throat, see e.g. [4, 6].
Detailed information on the thin-shell formalism in general relativity can be found in, e.g. [5, 7, 8] and references therein. The thin shell approach has had wide applications in general relativity and has been the subject of many studies. An initial study of the stability of the thin shell Schwarzschild wormhole about a static solution can be found in [3]. Since then many others have utilized similar techniques with varying spacetime metrics and conditions. Some studies have introduced a cosmological constant through the Schwarzschild - de Sitter, and Schwarzschild - anti-de Sitter spacetimes, e.g. [6, 9, 10, 11, 12, 13]. Other studies have utilized charged wormholes (Reissner–Nordström) [14, 15, 16, 17, 18, 19], rotating wormholes [20, 21], or Bardeen de-sitter wormholes [22]. Furthermore, the stability of wormholes has also been considered for modified theories of gravity such as $F(R)$ [17, 23], Einstein-Guass-Bonnet [24, 25, 26], and Hadara (Conformal Killing) gravity [22]. There has also been a study on the stability of various wormhole types when constrained by current cosmological observations [9]. Surprisingly, fewer studies [27, 28, 29, 30] have utilized the Friedmann-Lemaitre-Robertson-Walker metric, which we include in our paper along with other spacetimes.
In this paper, we explore and systematize the study of the stability conditions of spherically symmetric thin shell spacetime junctions considering linearized perturbations about a static equilibrium. It is demonstrated that stability exists in the casual region (where the perturbation sound speed is real and sub-luminal) for black holes. However, for all wormhole constructions that are explored, stability is found to be present only far outside this region.
The structure of the paper is described as follows. In Section II, we begin by establishing and summarizing the thin shell formalism, using the equation of motion of the radius of the throat to derive stability conditions. In Section III, the stability conditions are applied to wormhole and black hole junctions composed of Schwarzschild, Schwarzschild - de Sitter, and Schwarzschild anti-de Sitter spacetimes. A taxonomy of major categories is defined, and mathematical conditions related to the geometry of the stability regions for each category are derived (this portion will build and expand [6]. Next, in Section IV we consider the construction of wormholes and black holes using the Friedmann–Lemaître–Robertson–Walker (FLRW) metric and a Schwarzschild or Schwarzschild - (anti-) de Sitter metric. We derive the extrinsic curvature and use thin shell formalism to give and categorize the stability conditions of these junctions. The taxonomic conditions from the earlier section are generalized to apply to this latter section. In Section V, we present our analysis with plots and discussion. And finally, in Section VI a summary and concluding remarks are provided.
II Formalism
(a) Black hole
(b) Wormhole
Figure 1: Schwarzschild – Schwarzschild Junction: For the black hole, $M\neq 0$ everywhere and there is no stability flip. Stability region is always underneath the surface, partially intersecting the Causal Region (i.e., $10<P^{\prime}/\sigma^{\prime}<$ 1). For the wormhole, there is an asymptote and stability region is above surface A and below surface B.
(a) Black hole
(b) Wormhole
Figure 2: Schwarzschild - de Sitter – Schwarzschild Junction: Plot has been expanded out to show horizons. The de Sitter Horizon is visible, as expected for a junction containing a $\Lambda>0$  [^2] The horizon occurs at $R/m^{+}\approx 54$ which is close to $\alpha_{\rm dS}\approx 54.77$ for $C=0.001$ – see text. For the black hole, stability is defined by the vertical surface bifurcating the plot where $M=0$. Stability regions are below surface A and above surface B. A Schwarzschild - de Sitter – Schwarzschild junction implies $[C]=-0.001$ which does not fulfill the asymptote condition for any beta. Stability regions for the wormhole are below surface B and above surface A which is too small to be visible in this plot.
(a) Black hole
(b) Wormhole
Figure 3: Schwarzschild Anti-de Sitter – Schwarzschild Junction: Black hole has an asymptote as $[C]=0.001$ which does fulfill the asymptote condition. Stability regions for black hole below A and above B. Though not shown, wormhole at high $R$ is very similar to Schwarzschild – Schwarzschild and does not possess a de Sitter Horizon. Stability regions are above A and below B.
(a) Black hole
(b) Wormhole
Figure 4: Schwarzschild – Schwarzschild - de Sitter Junction: Black hole is similar to Schwarzschild - anti-de Sitter – Schwarzschild (FIG. 3) with a similar asymptote, though in this case the de Sitter Horizon is also present. Again $[C]=0.001$ in this case, fulfilling the asymptote condition. Stability regions for black hole below A and above B. Wormhole case at large $R$ is similar to Schwarzschild - de Sitter – Schwarzschild (FIG. 2). Stability regions are above A and below B.
(a) Black hole
(b) Wormhole
Figure 5: Schwarzschild – Schwarzschild - anti-de Sitter Junction: Black hole is similar to Schwarzschild - de Sitter – Schwarzschild (figure 2) but does not de Sitter Horizon. $[C]=-0.001$ and there is no asymptote. Stability regions are above A and below B.
We consider two spacetimes $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$ defined by metrics $g_{\alpha\beta}^{+}$ and $g_{\alpha\beta}^{-}$. We define two hypersurfaces within each spacetime as $\Sigma^{+}$ and $\Sigma^{-}$ with intrinsic metrics of $g_{ij}^{+}$ and $g_{ij}^{-}$, respectively. $x_{\pm}^{\gamma}$ refers to the coordinates in $g_{\alpha\beta}^{\pm}$ and $\xi_{\pm}^{c}$ refers to the coordinates in $g_{ij}^{\pm}$. The parametric equation of the surface takes the form $F(x^{\alpha}(\xi^{a}))=0$ [4].
Throughout this work, we shall define $[A]\equiv A^{+}-A^{-}$ and $\bar{A}\equiv\frac{1}{2}(A^{+}+A^{-})$ for some quantity $A$. The first Darmois condition for joining a portion of $\mathcal{M}^{+}$ to a portion of $\mathcal{M}^{-}$ is (see, e.g., [4])
$[g_{ij}]=0.$ (1)
This implies $g_{ij}^{+}=g_{ij}^{-}=g_{ij}$ and $\Sigma^{+}=\Sigma^{-}=\Sigma$.
(a) Black hole
(b) Wormhole
Figure 6: Schwarzschild - de Sitter – Schwarzschild - de Sitter Junction: Wormhole case at large $R$ is similar to Schwarzschild - de Sitter – Schwarzschild (figure 2). Stability regions are above A and below B. Black hole still possesses asymptote as $[C]=0$ fulfills asymptote condition in equation (27) if there exists a $\Lambda>0$. The location and shape of the asymptote differs from other cases and occurs at much higher $R$.
The coordinates $(\tau,\theta,\phi)$ are indicated by $\xi^{c}$ and $(t,r,\theta,\phi)$ are indicated by $x^{\gamma}$. $\dot{A}$ is defined as the derivative of $A$ with respect to proper time $\tau$. We adopt $c=G=1$ units throughout this work.
Per the thin shell approach, we let the throat of the wormhole be infinitesimally small and let each manifold have a boundary at the surface. In the case of a time-like spherically symmetric surface of dynamic radius $R(\tau)$ the surface line element can be written as [6]
$ds^{2}_{\Sigma}=-d\tau^{2}+R^{2}(\tau)d\Omega^{2},$ (2)
where the surface is defined by $r=R(\tau)$ and is parameterized by the function $F(r)=r-R(\tau)=0$.
This boundary will cause a discontinuity in the extrinsic curvature (second fundamental form) of the union of $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$. The stress-energy tensor ($S_{ij}$) for this boundary can be calculated using the Lanczos equation, which is given by [4]
$S_{ij}=-\frac{1}{8\pi}([K_{ij}]-g_{ij}[K^{i}_{i}]),$ (3)
where $K_{ij}$ is the extrinsic curvature and is given by
$K_{ij}=-n_{\gamma}\bigg{(}\frac{\partial^{2}x^{\gamma}}{\partial\xi^{i}\partial\xi^{j}}+\Gamma^{\gamma}_{\alpha\beta}\frac{\partial x^{\alpha}}{\partial\xi^{i}}\frac{\partial x^{\beta}}{\partial\xi^{j}}\bigg{)}.$ (4)
Note that $n_{\gamma}$ is the unit 4-normal to the surface $\Sigma$ in manifold $\mathcal{M}$ and is expressed as [4]
$n_{\gamma}=\pm\frac{1}{\big{(}\big{|}g^{\alpha\beta}\frac{\partial F}{\partial x^{\alpha}}\frac{\partial F}{\partial x^{\beta}}\big{|}\big{)}^{1/2}}\frac{\partial F}{\partial x^{\gamma}},$ (5)
where the sign of $n_{\gamma}$ depends on the direction of the normal vector.
We will treat $S_{ij}$ analogously to a perfect 4-fluid with $S_{ij}={\rm diag}(-\sigma,P,P)$ [3]. It can be shown that the energy density is
$\sigma(\xi^{a})=-S_{\tau}^{\tau}=-\frac{1}{4\pi}\big{[}K_{\theta}^{\theta}\big{]}.$ (6)
We define the mass of the thin shell as
$M=4\pi R^{2}\sigma=-\big{[}K_{\theta\theta}\big{]}.$ (7)
If equation (1) holds and $[K_{\theta\theta}]=0$ then we refer to $\Sigma$ as a boundary surface, if $[K_{\theta\theta}]\neq 0$ then $\Sigma$ is a thin shell.
22footnotetext: The horizon occurs at $R/m^{+}\approx 54$ which is close to $\alpha_{\rm dS}\approx 54.77$ for $C=0.001$ – see text
In Section III, we will consider a junction between two spacetimes of the form,
$ds_{\pm}^{2}=-\bigg{(}1-\frac{2\mu^{\pm}(r)}{r}\bigg{)}dt^{2}+\frac{dr^{2}}{1-\frac{2\mu^{\pm}(r)}{r}}+r^{2}d\Omega^{2},$ (8)
where $\mu^{\pm}(r)$ represents the effective mass contained within the thin shell and is defined as [4]
$\mu^{\pm}(r)=\frac{1}{2}(g^{\pm}_{\theta\theta})^{\frac{3}{2}}R_{\theta\phi}^{\quad\theta\phi},$ (9)
where metric tensor element $g_{\theta\theta}$ and Riemann tensor element $R_{\theta\phi}^{\quad\theta\phi}$ are computed on $\mathcal{M}$ (not on $\Sigma$).
Using equation (7) we get
$M(R)=wR\sqrt{1-\frac{2\mu^{-}}{R}+\dot{R}}-R\sqrt{1-\frac{2\mu^{+}}{R}+\dot{R}},$ (10)
where $w$ is determined by the direction of the normal vectors. If the vectors point in the opposite directions $w=-1$ and the junction is referred to as a wormhole. If the vectors point in the same direction $w=1$ and the junction is referred to as a black hole.
Rearranging equation (10) gives the equation of motion [4]
$\dot{R}^{2}=\bigg{(}\frac{[\mu]}{M}\bigg{)}^{2}+\frac{2\bar{\mu}}{R}+\bigg{(}\frac{M}{2R}\bigg{)}^{2}-1.$ (11)
If $\mu^{\pm}(R)$ is defined uniquely for each $R$ we can define a potential $V(R)=-\dot{R}^{2}$ [6]. We can expand this potential to second order about a static solution at $R_{0}=\rm const$. [3]
$V(R)\approx V(R_{0})+V^{\prime}(R_{0})(R-R_{0})+\frac{1}{2}V^{\prime\prime}(R_{0})(R-R_{0})^{2}.$ (12)
For this static solution, $V(R_{0})=V^{\prime}(R_{0})=0$. A stable solution will be given by $V^{\prime\prime}(R_{0})>0$. It follows from the definition of $V(R)$ that the equilibrium condition $V^{\prime}(R)=0$ for a static solution is
$\bigg{(}\frac{M}{2R}\bigg{)}^{\prime}=-\frac{2R}{M}\bigg{(}\bigg{(}\frac{(\mu]}{R}\bigg{)}\bigg{(}\frac{[\mu]}{R}\bigg{)}^{\prime}+\bigg{(}\frac{\bar{\mu}}{R}\bigg{)}^{\prime}\bigg{)}\equiv\Gamma.$ (13)
The condition for a stable equilibrium ($V^{\prime\prime}(R)>0$) is
$\bigg{(}\frac{M}{2R}\bigg{)}\bigg{(}\frac{M}{2R}\bigg{)}^{\prime\prime}<\Psi-\Gamma^{2},$ (14)
where $\Psi=-\big{(}\frac{[\mu]}{M}\big{)}^{\prime 2}-\big{(}\frac{[\mu]}{M}\big{)}\big{(}\frac{[\mu]}{M}\big{)}^{\prime\prime}-\big{(}\frac{\bar{\mu}}{R}\big{)}^{\prime\prime}$ [6].
In general, the conservation identity must also be satisfied and is given by
$\nabla_{i}S^{i}_{j}=-\bigg{[}T_{\alpha}^{\beta}\frac{\partial x^{\alpha}}{\partial\xi^{j}}n_{\beta}\bigg{]},$ (15)
which yields,
$\dot{\sigma}=-2\frac{\dot{R}}{R}(\sigma+P)+\Xi$ (16)
where the flux term $\Xi$ is
$\Xi\equiv\bigg{[}T_{\alpha}^{\beta}\frac{\partial x^{\alpha}}{\partial\tau}n_{\beta}\bigg{]}.$ (17)
For any vacuum solution, flux term $\Xi=0$, and the conservation identity becomes
$\sigma^{\prime}=-\frac{2}{R}(\sigma+P),$ (18)
which can be rewritten as,
$\bigg{(}\frac{M}{2R}\bigg{)}^{\prime\prime}=\frac{\Upsilon}{2R^{3}}\bigg{(}1+2\frac{P^{\prime}}{\sigma^{\prime}}\bigg{)},$ (19)
where $\Upsilon=3M-(MR)^{\prime}$. Plugging this into the stability condition (equation 14) yields [6]
$\frac{P^{\prime}}{\sigma^{\prime}}$ $<$ $\frac{1}{2}(\Phi-1);\quad M\Upsilon>0,\rm and$ (20)
$\frac{P^{\prime}}{\sigma^{\prime}}$ $>$ $\frac{1}{2}(\Phi-1);\quad M\Upsilon<0,$ (21)
where $\Phi=\frac{4R^{4}}{M\Upsilon}(\Psi-\Gamma^{2})$.
However, it is not necessary to compute $\Phi$. When using the condition $V^{\prime}=0$ one can show that the stability conditions become
$\frac{P^{\prime}}{\sigma^{\prime}}$ $<$ $\frac{R^{3}}{\Upsilon}\bigg{(}\frac{M}{2R}\bigg{)}^{\prime\prime}-\frac{1}{2};\quad M\Upsilon>0\rm;$ (22)
$\frac{P^{\prime}}{\sigma^{\prime}}$ $>$ $\frac{R^{3}}{\Upsilon}\bigg{(}\frac{M}{2R}\bigg{)}^{\prime\prime}-\frac{1}{2};\quad M\Upsilon<0.$ (23)
Note that the stability regions will lie above and below the surface mapped out by the transparency condition (equation 19).
Stability within the region defined by $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1 is of particular interest as it is the causal region. $P^{\prime}/\sigma^{\prime}$ corresponds to the square of a sound speed of perturbations. Thus, for physical solutions it seems natural to restrict this sound speed to real, sub-luminal values, where it is causal. The legitimacy of this assumption is discussed in the conclusion.
III Schwarzschild and (Anti-) de Sitter spacetimes
We begin our stability analysis by considering Schwarzschild wormhole and black hole solutions. We will also consider the effects of a positive and negative cosmological constant corresponding to Schwarzschild - de Sitter and Schwarzschild - anti-de Sitter spacetime respectively.
A Schwarzschild spacetime is one where $\mu^{\pm}(R)=m^{\pm}$ where $m^{\pm}$ is the Schwarzschild mass. A Schwarzschild - de Sitter spacetime is one where $\mu^{\pm}(R)=m^{\pm}+\frac{\Lambda^{\pm}}{6}R^{3}$ and $\Lambda^{\pm}>0$. And a Schwarzschild - anti-de Sitter spacetime is like the above but with $\Lambda^{\pm}<0$.
Applying the condition of a static solution with constant $R$, we can set $V=-\dot{R}=0$.
We now have $M$ given by
$M=w\sqrt{1-\frac{2m^{-}}{R}-\frac{\Lambda^{-}}{3}R^{2}}-\sqrt{1-\frac{2m^{+}}{R}-\frac{\Lambda^{+}}{3}R^{2}}.$ (24)
From these three possibilities we get 18 junction combinations (9 wormholes and 9 black holes). A taxonomy of the qualitative properties of the stability conditions, (equations 22, and 23) can be created by looking at two main features. The first is the existence of an asymptote at $\Upsilon=0$. The second is the limiting behavior as $R\rightarrow\infty$, whether or not the surface diverges at some finite value. Finally, it is also worth noting the presence of a stability flip (the stability region shifting from above the surface to below or vice versa) caused by $M$ or $\Upsilon$ switching signs.
Throughout the rest of this paper where relevant, we will only consider cases where $m^{-}<m^{+}$, which is reflected in our axis limits of our plots $10<m^{-}/m^{+}<$ 1 [6] (in general $\mu^{-}$ and $\mu^{+}$ can be swapped without consequence to the stability regions).
First, it is important to note the region for which our analysis cannot yield results. When $R<2\mu_{\rm max}$ the radius of our thin shell is within the event horizon of one of the Schwarzschild space times and due to our use of Schwarzschild coordinates, we cannot analyze stability. Similarly, when at least one Schwarzschild - de Sitter spacetime is used, a cosmological horizon known as the de Sitter horizon is present at high $R$. This causes the surface to diverge at finite $R$. Since $R$ is large at this horizon we can estimate its value as $R_{\rm dS}\approx\sqrt{3/\Lambda_{\rm max}}$ where $\Lambda_{\rm max}$ is the greatest $\Lambda$ involved in the junction. Beyond the horizon, we reach another region where our stability analysis breaks down.
An asymptote between these two horizons will occur if $\Upsilon=0$ [6]. A wormhole with any combination of the above spacetimes will always contain such an asymptote. A black hole may or may not contain an asymptote based on the following.
It can be seen from plotting that for small $R$ (near the event horizon) $\Upsilon$ and $M$ both have the same sign. With the exception of Schwarzschild – Schwarzschild (FIG. 1), after a certain point (before $\Upsilon=0$ or $M=0$) if $\Upsilon$ monotonically increases, $M$ monotonically decreases or vice versa. Thus, only one of these quantities can be equal to zero for a certain junction. If $M=0$ anywhere then $\Upsilon\neq 0$ and there is no asymptote, though a stability flip will still exist at $M=0$. This lets the simpler condition $M=0$ become an indicator of the nonexistence of an asymptote.
$M=0$ implies $\mu^{+}=\mu^{-}$ or
$\frac{6[m]}{R^{3}}=-[\Lambda].$ (25)
Since $R$ must be positive, if $[m]$ and $[\Lambda]$ have the same sign, the condition can never be satisfied for any $R$, and so an asymptote must exist. This gives an asymptote existence condition of
$\frac{[\Lambda]}{[m]}>0.$ (26)
For a junction with at least one $\Lambda>0$ there is an upper limit to $R$. If the $R$ value which satisfies the horizon condition (equation 25) is greater than $R_{\rm dS}$ then the condition for asymptote non-existence cannot be met and an asymptote will be present. Taking $R_{\rm dS}\approx\sqrt{3/\Lambda_{\rm max}}$ gives an asymptote existence condition of
$\frac{[\Lambda]}{[m]}>\frac{-2\Lambda_{\rm max}^{\frac{3}{2}}}{\sqrt{3}}.$ (27)
For both $\Lambda\leq 0$ this restriction does not exist and only equation (26) applies.
For the Schwarzschild – Schwarzschild case (FIG. 1), $M$ and $\Upsilon$ approach zero for large $R$, but never reach it. Thus, there is neither an asymptote or stability flip in this simple case.
It is also apparent from plotting that $M\Upsilon$ will monotonically decrease for black holes and monotonically increase for wormholes. This behavior dictates the location of stability regions, whether above or below the surface.
Table 1: Notable Features of Black hole Parameter Spaces.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability regions Stable in causal region Figure
Sch – Sch No Convergent Below Yes 1
Sch – Sch-deSit Yes Divergent Below A, Above B Yes 4
Sch – Sch-Anti deSit No Convergent Below A, Above B Yes 5
Sch-deSit – Sch No Divergent Below A, Above B Yes 2
Sch-deSit – Sch-deSit Yes Divergent Below A, Above B Yes 6
Sch-deSit – Sch-Anti deSit No Divergent Below A, Above B Yes -
Sch-Anti deSit – Sch Yes Convergent Below A, Above B Yes 3
Sch -Anti deSit – Sch-deSit Yes Divergent Below A, Above B Yes -
Sch-Anti deSit – Sch-Anti deSit No Convergent Below Yes 7
FLRW k=0 – Sch No Convergent Below A, Above B Yes 8
FLRW k=0 – Sch-deSit No Convergent Below A, Above B Yes 9
FLRW k=0 – Sch-Anti deSit No Convergent Below A, Above B Yes 10
FLRW k=+1 – Sch No Divergent Below A, Above B Yes 11
FLRW k=-1 – Sch No Convergent Below A, Above B Yes 12
(a) Black hole
(b) Wormhole
Figure 7: Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter Junction: Both wormhole and black hole are very similar to Schwarzschild – Schwarzschild junction (figure 1). The wormhole parameter surface takes the usual form (stability above A and below B). The black hole has no asymtote and approaches 0 as $R\rightarrow\infty$. $M\neq 0$ for any $\beta$ and there is no stability flip. Stability region is always below the surface.
Table 2: Notable Features of Wormhole parameter Spaces.
Type Asymptote at $\Upsilon=0$ Limit as $R\rightarrow\infty$ Stability Regions Stable in Causal Region Figure
Sch – Sch Yes Convergent Above A, Below B No 1
Sch – Sch-deSit Yes Divergent Above A, Below B No 4
Sch – Sch-Anti deSit Yes Convergent Above A, Below B No 5
Sch-deSit – Sch Yes Divergent Above A, Below B No 4
Sch-deSit – Sch-deSit Yes Divergent Above A, Below B No 6
Sch-deSit – Sch-Anti deSit Yes Divergent Above A, Below B No -
Sch-Anti deSit – Sch Yes Convergent Above A, Below B No 3
Sch -Anti deSit – Sch-deSit Yes Divergent Above A, Below B No -
Sch-Anti deSit – Sch-Anti deSit Yes Convergent Above A, Below B No 7
FLRW k=0 – Sch Yes Convergent Above A, Below B No 8
FLRW k=0 – Sch-deSit Yes Convergent Above A, Below B No 9
FLRW k=0 – Sch-Anti deSit Yes Convergent Above A, Below B No 10
FLRW k=+1—FLRW Yes Divergent Above A, Below B No 11
FLRW k=-1—FLRW Yes Convergent Above A, Below B No 12
IV FLRW spacetime
In this section, we turn our attention to a junction between a spacetime defined by the Friedmann-Lemaître-Robertson-Walker (FLRW) metric and a Schwarzschild or Schwarzschild - (anti-) de Sitter spacetime. The FLRW line element is defined as
$ds^{2}=-c^{2}dt^{2}+a^{2}(t)\bigg{(}\frac{dr^{2}}{1-kr^{2}}+r^{2}d\Omega^{2}\bigg{)},$ (28)
where $a(t)$ is the scale factor.
We denote the FLRW spacetime as $\mathcal{M}^{-}$ and the Schwarzschild spacetime as $\mathcal{M}^{+}$.
Unlike the previous stationary spacetimes, the FLRW spacetime is expanding. Here we follow the precedent established in the Swiss Cheese Cosmological Model, which considers spherical regions of Schwarzschild spacetime matched to an FLRW background [31]. In these cases, it is typical to define a hypersurface that is expanding at the same rate as the FLRW so that in the FLRW frame it is stationary apart from the evolution of the hypersurface radius defined in equation (11). The surface in the FLRW frame will be defined by $r_{\rm f}=R(\tau)$. The static solution solved for here will be one for which $\dot{R}=0$.
In the Schwarzschild (or Schwarzschild - (Anti) de Sitter) frame, we define the surface as $r_{\rm s}=a(t(\tau))R(\tau)\equiv\chi(\tau)$. This allows our junction to satisfy the first Darmois condition (equation 1) as in the Swiss Cheese Model. Unlike the Swiss Cheese Model, however, we do not need to satisfy the second Darmois condition as we are using thin shells which cause a discontinuity in extrinsic curvature [30].
The line element of our new hypersurface becomes
$ds^{2}_{\Sigma}=-d\tau^{2}+a^{2}(t)R^{2}(\tau)d\Omega^{2}=-d\tau^{2}+\chi^{2}(\tau)d\Omega^{2}.$ (29)
The introduction of an expanding hypersurface invalidates many of the assumptions made in previous sections. Firstly, flux term $\Xi$ is now non-zero and can be expressed as
$\Xi=\pm\Bigg{[}\dot{r}(\rho+p){\bigg{(}\bigg{|}g^{\alpha\beta}\frac{\partial F}{\partial x^{\alpha}}\frac{\partial F}{\partial x^{\beta}}\bigg{|}\bigg{)}^{-1/2}}\Bigg{]}$ (30)
where $\rho$ and $p$ are the energy density and pressure of a perfect 4-fluid. When considering an FLRW – Schwarzschild junction, the flux term becomes
$\Xi=\mp\frac{\rho_{m}a\dot{R}}{\sqrt{1-kR^{2}-\big{(}\frac{dR}{dt}\big{)}^{2}}},$ (31)
where $\rho_{m}$ is the matter energy density.
Taking $\dot{\chi}=\dot{a}R+\dot{R}a$, the conservation identity now yields
$\dot{\sigma}=-2\bigg{(}\frac{\dot{R}}{R}+\frac{\dot{a}}{a}\bigg{)}(\sigma+P)+\frac{\rho_{m}\dot{R}}{\sqrt{1-kR^{2}}}.$ (32)
For a static solution $\dot{R}=0$, the identity becomes
$\dot{\sigma}=-2\frac{\dot{a}}{a}(\sigma+P),$ (33)
and the flux term does not affect the equation.
Here we break away from using prime notation to express derivatives with respect to $R$ and instead let $A^{\prime}\equiv\frac{\partial A}{\partial\chi}$. Using this notation, equation (33) becomes
$\sigma^{\prime}=-\frac{2}{\chi}(\sigma+P).$ (34)
Note that $\sigma$ is a function of $\chi(\tau)$.
Here we will let $M=-[K_{\theta\theta}]=4\pi\chi^{2}\sigma$. Using this, the continuity equation can finally be expressed as
$\frac{P^{\prime}}{\sigma^{\prime}}=\frac{\dot{P}}{\dot{\sigma}}=\frac{\chi^{3}}{\upsilon}\bigg{(}\frac{M}{2\chi}\bigg{)}^{\prime\prime}-\frac{1}{2},$ (35)
where $\upsilon\equiv 3M-(M\chi)^{\prime}$.
Now, we can derive the mass of the thin shell, $M$, using (equation 7).
For ease of computation, we define $f(r)=1-\frac{2\mu(r)}{r}$ for the Schwarzschild and Schwarzschild - (anti-) de Sitter spacetime.
Calculating the Riemann tensor for FLRW and using equation (9), we get
$\mu(R)=\frac{\chi^{3}}{2}\bigg{(}H^{2}+\frac{k}{a^{2}}\bigg{)},$ (36)
where $H$ is the Hubble parameter defined as $H=\frac{1}{a}\frac{da}{dt}$ in the FLRW frame.
(a) Black hole
(b) Wormhole
Figure 8: FLRW k=0 – Schwarzschild Junction: The black hole and wormhole graphs are similar to previous examples though does not posses a de Sitter Horizon. It is worth noting that the radius of the event horizon decreases rapidly at low $a$ due to a high value of $H^{2}$. This is much easier to see in the wormhole graph. Stability regions for the black hole are once again below surface A, above surface B. For wormhole, stability regions are above surface A and below surface B.
(a) Black hole
(b) Wormhole
Figure 9: FLRW k=0 – Schwarzschild - de Sitter Junction: Here $[C]=0$ causing the $M=0$ to be linear for the black hole. Outside of this, plots are very similiar to figure (8). For black hole, stability regions are below A and above B. For wormhole stability regions are above A and below B.
The normal vector for Schwarzschild becomes
$n_{\gamma}=\pm\bigg{(}f-\frac{1}{f}\bigg{(}\frac{d\chi}{dt}\bigg{)}^{2}\bigg{)}^{-1/2}\bigg{(}-\frac{d\chi}{dt},1,0,0\bigg{)},$ (37)
and for the FLRW
$n_{\gamma}=\pm\bigg{(}\frac{1-kR^{2}}{a^{2}}-\bigg{(}\frac{dR}{dt}\bigg{)}^{2}\bigg{)}^{-1/2}\bigg{(}-\frac{dR}{dt},1,0,0\bigg{)}.$ (38)
The $\theta\theta$ element of the second fundamental form for both simplifies to
$K_{\theta\theta}=-n_{\gamma}\Gamma^{\gamma}_{\theta\theta}.$ (39)
For Schwarzschild,
$K_{\theta\theta}^{+}=\pm\chi f^{3/2}\bigg{(}f^{2}-\bigg{(}\frac{d\chi}{dt}\bigg{)}^{2}\bigg{)}^{-1/2}.$ (40)
For FLRW, $K_{\theta\theta}^{-}$ yields a complicated form. We simplify this by taking $K_{\theta\theta}$ at the static solution $\dot{R}=\frac{dR}{dt}\frac{dt}{d\tau}=0\implies\frac{dR}{dt}=0$ if $\frac{dt}{d\tau}\neq 0$.
Thus,
$K_{\theta\theta}^{-}=\pm\chi\sqrt{1-kR^{2}}.$ (41)
From the first Darmois condition (equation 1) and using the definition of the induced metric we can find an equation for $\frac{dt}{d\tau}$ for both spacetimes.
For Schwarzschild:
$\bigg{(}\frac{dt}{d\tau}\bigg{)}^{2}=\frac{1}{f^{2}}(f+\dot{\chi}^{2}),$ (42)
and for FLRW:
$\bigg{(}\frac{dt}{d\tau}\bigg{)}^{2}=\frac{a^{2}}{1-kR^{2}}\dot{R}^{2}+1.$ (43)
Plugging into the Schwarzschild $K_{\theta\theta}$ yields
$K_{\theta\theta}^{+}$ $=$ $\chi\sqrt{f+\dot{\chi}},$ (44)
$=$ $\chi\sqrt{1-\frac{2\mu(\chi)}{\chi}+\dot{\chi}^{2}}.$
For a static solution, in the FLRW frame
$aH=\frac{da}{dt}=\dot{a}\frac{dt}{d\tau}=\dot{a}.$ (45)
This implies
$\dot{\chi}=\dot{a}R=H\chi.$ (46)
Utilizing the above and (equation 36) we can express the $K_{\theta\theta}$ for the FLRW spacetime in the same manner,
$K_{\theta\theta}^{-}=\chi\sqrt{1-\frac{2\mu(\chi)}{\chi}+\dot{\chi}^{2}}.$ (47)
Using (equation 7) we can derive a new equation of motion,
$\dot{\chi}^{2}=\bigg{(}\frac{[\mu]}{M}\bigg{)}^{2}+\frac{2\bar{\mu}}{\chi}+\bigg{(}\frac{M}{2\chi}\bigg{)}^{2}-1=-V(\chi(R)),$ (48)
which is in the same form as before.
Though, at the static solution $V=H\chi\neq 0$ generally. $V^{\prime}$ may be be chosen to be 00 for an equilibrium point. Thus, $V^{\prime\prime}>0$ will again yield stable equilibria. Combining this with the transparency condition (equation 35) and the fact that $V^{\prime}=0$ at this point, gives stability conditions of
$\frac{P^{\prime}}{\sigma^{\prime}}$ $<$ $\frac{\chi^{3}}{\upsilon}\bigg{(}\frac{M}{2\chi}\bigg{)}^{\prime\prime}-\frac{1}{2};\quad M\upsilon>0\rm,$ (49)
$\frac{P^{\prime}}{\sigma^{\prime}}$ $>$ $\frac{\chi^{3}}{\upsilon}\bigg{(}\frac{M}{2\chi}\bigg{)}^{\prime\prime}-\frac{1}{2};\quad M\upsilon<0.$ (50)
Finally, we can express $M$ in a static case where $\dot{\chi}=H\chi$ as
$M=w\chi\sqrt{1-k\frac{\chi^{2}}{a^{2}}}-\chi\sqrt{1-\frac{2m}{\chi}-\bigg{(}\frac{\Lambda^{+}}{3}-H^{2}\bigg{)}\chi^{2}}.$ (51)
It is now clear that each new $K_{\theta\theta}$ takes on the same form as before but now has an additional positive term $\dot{\chi}=H\chi$ that incorporates the expansion. This positive term acts like a negative mass. For example, in the Schwarzschild spacetime, the expansion term allows an event horizon radius smaller than the traditional Schwarzschild radius.
$M$ may also be recast into a form analogous to equation (24) as
$M=w\chi\sqrt{1-\frac{2m^{-}}{\chi}-\frac{\varepsilon^{-}\chi^{2}}{3}}-\chi\sqrt{1-\frac{2m^{+}}{\chi}-\frac{\varepsilon^{+}\chi^{2}}{3}}.$ (52)
Instead of using $\Lambda$, we define $\varepsilon/3$ to be the coefficient of the $\chi^{2}$ term in each $K_{\theta\theta}$.
(a) Black hole
(b) Wormhole
Figure 10: FLRW k=0 – Schwarzschild - anti-de Sitter Junction: Very similar to figure (8). $M=0$ plane is shifted to lower $\chi$ values. Stability regions for black hole are below A, above B. Stability regions for wormhole are above A, below B.
We may now use the same asymptote conditions as in Section III just substituting $\varepsilon$ for $\Lambda$.
The first Friedmann Equation is
$H^{2}=\frac{8\pi}{3}\rho-\frac{k}{a^{2}}+\frac{\Lambda^{-}}{3}.$ (53)
For Schwarzschild or Schwarzschild - (anti-) de Sitter, we can define $\varepsilon$ as
$\varepsilon^{+}=\Lambda^{+}-3H^{2}=\frac{3k}{a^{2}}-8\pi\rho+[\Lambda],$ (54)
where $\Lambda^{+}$ is the cosmological constant of the Schwarzschild - (anti-) de Sitter and $\Lambda^{-}$ is the cosmological constant of the FLRW. For Schwarzschild, $\Lambda^{+}=0$.
For FLRW $\varepsilon$ is simply
$\varepsilon^{-}=\frac{3k}{a^{2}}.$ (55)
For all Schwarzschild spacetimes, $m^{+}=m$, where $m$ is the Schwarzschild mass and is constant. For FLRW, $m^{-}=0$.
The condition for the existence of a de Sitter horizon becomes $\varepsilon^{+}>0$ or $\varepsilon^{-}>0$ which can be expressed as
$[\Lambda]>8\pi\rho-\frac{3k}{a^{2}},$ (56)
or
$3k>0\implies k=+1,$ (57)
since $k$ must be either -1, 0, or +1.
Lastly, the asymptotic conditions (equations 26, 27) are re-expressed for $\upsilon=0$ as
$\frac{[\varepsilon]}{[m]}>0,$ (58)
$\frac{[\varepsilon]}{[m]}>\frac{-2\varepsilon_{\rm max}^{\frac{3}{2}}}{\sqrt{3}}.$ (59)
(a) Black hole
(b) Wormhole
Figure 11: FLRW k=+1 – Schwarzschild Junction: For $k\neq 0$ we choose $m=0.02$. Note that $k>0$ gives $H^{2}<0$ for some values of $a(t)$. These non-physical solutions have been removed from the plots. A horizon caused by the positive curvature exists at high $\chi$ for both black hole and wormhole, though is more visible in the black hole. This horizon follows along the plane defined by $\chi=\frac{a}{m}$. At larger $m$ values the asymptote condition can be fulfilled but commonly this occurs only when $H^{2}<0$. Stability regions are below A and above B for black hole and are above A and below B for wormhole.
V Results
In the following, we will turn our attention to the plotting of regions of stability in three-dimensional parameter space. Overall qualitative features of the observed stability regions will be discussed.
V.1 Plotting
When utilizing the results of the previous sections, it is helpful to make a coordinate transformation for easier plotting. We consider $\alpha=R/m^{+}$ $\beta=m^{-}/m^{+}$, and define $C^{\pm}=\Lambda^{\pm}m^{+2}$.
For Schwarzschild or Schwarzschild - (anti-) de Sitter, we have
$M=wm^{+}\alpha\sqrt{1-2\frac{\beta}{\alpha}-\frac{C^{-}\alpha^{2}}{3}}-m^{+}\alpha\sqrt{1-2\frac{1}{\alpha}-\frac{C^{+}\alpha^{2}}{3}}.$ (60)
We can now view the parameter space by plotting $P^{\prime}/\sigma^{\prime}$ as a function of $\alpha$ and $\beta$. Since $m^{+}$ is not a function of $R$, when plugging into equations (22) and (23) it will cancel and have no effect on the final stability surface. Following a similar approach as [6], we shall let $|C^{\pm}|=|\Lambda^{\pm}m^{+2}|\sim 10^{-3}$. This assumption gives a high value for this product, but is useful for qualitative plotting.
Applying the substitutions to the asymptotic conditions given by equations (26) and (27) yields
$[C]>0,$ (61)
since $01-\beta>$ 0, and
$[C]>\frac{-2C_{\rm max}^{\frac{3}{2}}}{\sqrt{3}}(1-\beta).$ (62)
It can also be seen that $\alpha_{\rm\rm dS}\approx\sqrt{3/C_{\rm max}}$.
For the junctions between FLRW and Schwarzschild, we consider $\alpha=\chi/m$ and $C^{\pm}=\Lambda^{\pm}m^{2}$. Note that $\beta=m^{-}/m=0$ in this case as $m^{-}=0$. Thus, for plotting, we will use the $\alpha$ and $a(t)$ axes.
We choose $C^{+}=\pm 10^{-3}$ for Schwarzschild - de Sitter or Schwarzschild - anti-de Sitter and $C^{-}=+10^{-3}$ always. Thus, $[C]\leq 0$.
For the FLRW – Schwarzschild junctions, using the continuity equation of the FLRW spacetime, the evolution of $\rho$ can be expressed as $\left(\frac{\rho^{0}}{a^{3}}\right)$ where $\rho^{0}$ is a constant. We define $D\equiv 8\pi m^{2}\rho^{0}$ to keep dimensional consistency with $C$. Using our assumption for the magnitude of $C$ and the value of $\rho_{\Lambda}$ we find that for the present universe ($a(t)=1$) $D/C=\rho^{0}/\rho_{\Lambda}=\Omega_{m}^{0}/\Omega_{\Lambda}\approx 0.3/0.7$. Thus, $D\sim 4.23\times 10^{-4}$. From this, $M$ is as follows,
$M$ $=$ $wm\alpha\sqrt{1-\frac{km^{2}}{a^{2}}\alpha^{2}}$ (63)
$-m\alpha\sqrt{1-\frac{2}{\alpha}-\bigg{(}\frac{3km^{2}}{a^{2}}-\frac{D}{a^{3}}+[C]\bigg{)}\frac{\alpha^{2}}{3}}.$
Again, since $m$ is constant with respect to $\chi$, it will cancel when plugged into equations (49) and (50). However, the contribution from curvature $k$ is affected by the $m^{2}$ factor which cannot be ignored. A value for this $m$ must be chosen when plotting $k\neq 0$.
(a) Black hole
(b) Wormhole
Figure 12: FLRW k=-1 – Schwarzschild Junction: There is no horizon present at high $\chi$ for neither the black hole or the wormhole. For high $\chi$ the value of the surface approaches 0 instead of $\sim-0.5$. Stability regions are below A and above B. Surface B of wormhole graph decreases parabolically at higher $\chi$ than plotted. Stability regions are above A and below B.
Applying the substitutions to the conditions for the existence of a de Sitter Horizon (equations 56 and 57) gives
$[C]>\frac{D}{a^{3}}-\frac{3km^{2}}{a^{2}}.$ (64)
or
$k=+1.$ (65)
However, due the fact that $[C]\leq 0$, equation (64) is never satisfied when $k\neq+1$.
Applying the substitutions to the asymptotic conditions in equation (58) yields
$[C]>\frac{D}{a^{3}},$ (66)
which is never satisfied for $[C]\leq 0$.
When using equation (59) we assume $\varepsilon_{\rm max}=\varepsilon^{-}=3km^{2}$ as $D>0$ and $[C]<0$ cause $\varepsilon^{-}>\varepsilon^{+}$. This condition will only be relevant if $\varepsilon^{-}>0$ which implies the $k=+1$ case. So,
$[C]>\frac{D-6m^{3}}{a^{3}},$ (67)
which implies if $D-6m^{3}>0$ there can be no asymptote at any $a(t)$ for $[C]\leq 0$. Additionally, $\alpha_{\rm dS}\approx\frac{1}{m}$ when $k=+1$. Such a horizon is nonexistent for $k\neq+1$.
Plotted examples of parameter spaces for Schwarzschild and Schwarzschild - (anti-) de Sitter junctions are given in figures (1, 2, 3, 4, 5, 6, 7). Figures (8, 9, 10, 11, 12) detail parameter spaces of FLRW junctions with Schwarzschild and Schwarzschild - (anti-) de Sitter.
Throughout the plots, there are typically two separate regions of stability. Region A will refer to the stability region at lower $R$ or $\chi$ than the $\Upsilon=0$ asymptote or the $M=0$ stability flip. Region B shall refer to the higher $R$ or $\chi$ region [6].
V.2 Discussion
From the results seen in the figures throughout and summarized in Tables 1 and 2, we can see that the predictions made in Section III and IV are met. All (FLRW included) wormhole junctions possess an asymptote at $\Upsilon=0$. The shape and location of the asymptote can vary, but in all cases it divides the parameter space into two distinct stability regions. Stability region A is at lower $R$ or $\chi$ and is bound by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}>0$. Stability region B is at higher $R$ and is bound by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}<0$. In all cases observed these stability regions do not intersect with the Causal Region, $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1.
All (FLRW included) black hole junctions either possess an asymptote or an $M=0$ stability flip with the exception of the simple Schwarzschild – Schwarzschild and Schwarzschild - anti-de Sitter – Schwarzschild - anti-de Sitter. Of those that have some form of stability flip, (either due to $M=0$ or $\Upsilon=0$) there are two separate stability regions. Region A is at low $R$ or $\chi$ and is defined by a maximum $\frac{P^{\prime}}{\sigma^{\prime}}$. Region B is at higher R and is defined by a minimum $\frac{P^{\prime}}{\sigma^{\prime}}$. The maximum is above $\frac{P^{\prime}}{\sigma^{\prime}}=0$, causing an intersection of the stability regions with the Causal region. However, such an intersection does not exist for every combination of $\alpha$ and $\beta$ or $\alpha$ and $a(t)$.
In our plotting, since $m^{+}\geq m^{-}$, the swapping of spacetimes $\mathcal{M}^{+}$ and $\mathcal{M}^{-}$ has a noticeable effect on the stability regions. However, if one also changes the sign of $\Lambda$ in this swap, the qualitative features of the plot are preserved. For example, Schwarzschild Anti-de Sitter – Schwarzschild (figure 3) and Schwarzschild – Schwarzschild - de Sitter (figure 4) produce similar plots.
FLRW junctions act similar to junctions with only Schwarzschild, Schwarzschild - (anti-) de Sitter but lack asymptotes for black holes. Additionally, due to the positive expansion term $H^{2}\chi^{2}$, the event horizon approaches zero for low $a(t)$. A cosmological horizon is only permitted for $[C]>0$ (which is beyond the scope of this paper) or $k=+1$.
Each black hole and wormhole with and without an FLRW spacetime can be successfully grouped into the 4 categories, as can be seen in Tables 1 and 2.
VI Conclusion
We performed a stability analysis of thin shell wormholes and black holes constructed from combinations of Schwarzschild, Schwarzschild - de Sitter, Schwarzschild - anti-de Sitter, and Friedmann-Lemaître-Robertson-Walker (FLRW) spacetimes. We aimed at providing a taxonomy of these combinations while emphasizing some common landscape characteristics with the localization of stability and causality regions.
As previously stated, we found no parameter combinations which yield a stable wormhole when restricting $10\leq\frac{P^{\prime}}{\sigma^{\prime}}<$ 1. The boundaries of the stability regions consistently lie far outside the causal region, either at very high values or at negative ones. This conclusion has been echoed in previous works see e.g., [3, 6, 7, 32]. Our investigation of junctions containing the FLRW spacetime have yielded the same results.
One may interpret $P^{\prime}/\sigma^{\prime}$ as the square of the sound speed of perturbation, in which case it would be necessary to restrict its values to the causal region where the sound speed could not exceed the speed of light and be restricted to the reals. In this case, one would tend to deny the stability of the considered wormholes in the physical universe. However, as it was cautioned in [3, 7, 32], it may not be fully justified to entirely rule stability out. Indeed, the geometry of a wormhole already requires a violation of the null energy condition where the surface energy density $\sigma<0$ [3]. Such unusual conditions could allow for $P^{\prime}/\sigma^{\prime}$ to take on values outside of the causal range. It has been noted that negative $P^{\prime}/\sigma^{\prime}$ values can occur with the Casimir Effect and the False Vacuum [3]. Unfortunately, until a detailed model of exotic matter is formulated, it remains unclear if such values should be excluded from the discussion or not. It is also of note that possible stability in the causal region has been noted in rotating (BTZ metric) wormholes for sufficient angular momentum [20]. Thus, these wormholes do not require a relaxation of the $10<P^{\prime}/\sigma^{\prime}<$ 1 condition to be stable.
Finally, we have categorized the different junctions into taxonomic groups and depicted the mathematical conditions under which these groups are formed. While we found stability in the causal region only for black holes configurations, it remains possible that stable wormholes could exist, though only in extraordinary conditions which may not be allowed from the point of the view of semi-classical GR. This is still found to be the case even when FLRW spacetimes are used in the configurations.
Acknowledgements
We thank Francisco Lobo for useful comments on the manuscript. MI acknowledges that this material is based upon work supported in part by the Department of Energy, Office of Science, under Award Number DE-SC0022184 and also in part by the U.S. National Science Foundation under grant AST2327245.